OpenStudy (anonymous):

explain the difference between using the trigonometric ratios (sin,cos,tan) to solve for a missing angle in a right triangle versus using the reciprocal ratios (sec,csc,cot). you must use complete sentences and any evidence needed (such as example) to prove your point of view.

OpenStudy (anonymous):

medal will be given

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@jdoe0001 please help me

OpenStudy (anonymous):

medal will be given!!! help someone

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@Luigi0210 please help me

OpenStudy (anonymous):

@jdoe0001 please help me please

OpenStudy (anonymous):

I SERIOUSLY NEED HELP PLEASEEEEEEEEEEEE @Luigi0210

OpenStudy (anonymous):

@dan815 hi can you help me

OpenStudy (anonymous):

@wio can you please help me?

OpenStudy (anonymous):

medal will be given

OpenStudy (anonymous):

can someone help

OpenStudy (anonymous):

well \[ \sin = opposite/hypotenuse \]But \[ \csc = hypotenuse / opposite \]

OpenStudy (anonymous):

So you can use \[ hypotenuse \times \sin = opposite \]if you want to find the opposite side length and have hypotenuse. You can use \[ opposite \times \csc = hypotenuse \]If you want to find the hypotenuse and have the opposite side.

OpenStudy (anonymous):

but what is the difference of the two when you want to find an angle

OpenStudy (anonymous):

and in a right triangle

OpenStudy (anonymous):

in that case, it doesn't really matter.

OpenStudy (anonymous):

ok.. so the explaination that you gave me was to find an angle right?

OpenStudy (anonymous):

we dont have to find a side we have to find a missing angle.... @wio

OpenStudy (anonymous):

@jdoe0001 i need help

OpenStudy (anonymous):

@HelpBlahBlahBlah please help

OpenStudy (anonymous):

@thomaster please help

OpenStudy (anonymous):

@IMStuck hi can you pleae help me

OpenStudy (imstuck):

Oh, I will definitely try. Let me do some dwelling first...

OpenStudy (anonymous):

ok thankyou @IMStuck just please dont forget to come back

OpenStudy (imstuck):

When you use the trig functions as opposed to their cofunctions, there are no fractions involved. When you use cofunctions, such as sec, csc, and cot, you are involving fractions. For example, in this triangle

OpenStudy (imstuck):

So to find the sin of an angle using csc, you have to take the reciprocal of the result. It is the same with all the functions and their cofunctions. They are reciprocals.

OpenStudy (anonymous):

ok thank you so much!!!

OpenStudy (imstuck):

Does that help you in your endeavors to answer this correctly?

OpenStudy (anonymous):

yes it does unless you have something else to say

OpenStudy (anonymous):

wait does it turn into a negative @IMStuck

OpenStudy (imstuck):

no. They are not opposite reciprocals, only reciprocals.

OpenStudy (anonymous):

ok thank you so much

OpenStudy (imstuck):

You're very welcome!